OpenStudy (anonymous):

Solve -2(3-m)+15=6m-4(m-20) Please show all work

OpenStudy (anonymous):

-6+2m+15=6m-4m+80 2m-6m+4m=80-15+6 0=71 impossible no solution

OpenStudy (anonymous):

Do we have a name for a equation that don't have any solution ? (a formal name)

OpenStudy (anonymous):

if we do,i've forgotten it:(

OpenStudy (anonymous):

I believe it's inconsistent.